# White Dwarf Pollution; One Star or Two?

### 13 Mar 2024 v1

@@ Abstract
The accretion of tidally disrupted planetary bodies is the current consensus model for the presence of photospheric metals commonly detected in white dwarfs. While most dynamical studies have considered a single star and associated planetary instabilities, several investigations have instead considered the influence of widely-bound stellar companions as potential drivers of white dwarf pollution. This study examines the prevalence of wide binaries among polluted white dwarfs using Gaia DR3 astrometry, where three samples are investigated: 71 DAZ stars with metals detected in the ultraviolet using Hubble, and two groups of DZ stars identified via SDSS spectroscopy, comprised of 116 warmer and 101 cooler sources. Each sample was searched for spatially-resolved, co-moving companions, and compared to the same analysis of thousands of field white dwarfs within overlapping regions of the Gaia HR diagram. The wide binary fraction of the DAZ sample is per cent, and within of the corresponding field. However, the search yields wide binary fractions of less than 1.8 per cent for the two independent DZ star catalogues, which are each distinct from their fields by more than . Both sets of results support that pollution in white dwarfs is not the result of stellar companions, and the delivery of metals to white dwarf surfaces is caused by major planets. The discrepancy between the DAZ and DZ star wide binary fractions cannot be caused by white dwarf spectral evolution, suggesting these two populations may have distinct planetary architectures.
@@
Results to date have found the frequency of atmospheric pollution among single white dwarfs to be indistinct from those in binary systems (Wilson et al., 2019; Zuckerman, 2014) - findings that should be taken with caution as the sample statistics (12 and 38 binaries examined for pollution, respectively) are far from robust. A search for infrared excesses around an unbiased sample of 14 polluted white dwarfs in wide binaries, revealed none to have a significant flux or colour excess, while an independent assessment using Gaia DR2 astrometry further confirmed a duplicity fraction consistent with zero for the 40 known dusty white dwarfs with infrared excesses (Wilson et al., 2019). Previous efforts to identify binary companions or duplicity fractions among white dwarfs have been extremely limited by incomplete or inhomogeneous proper motion and photometric surveys (Farihi et al., 2005; Hollands et al., 2018b). Far superior constraints on the wide binary population of polluted white dwarfs can now be obtained with Gaia.
@@ 2 Sample Selection and Processing
There are three samples of metal-rich white dwarfs utilized to study the fraction of wide binary companions that may dynamically influence their respective planetary systems and pollution characteristics. The first sample is a subset of stars with photospheric metal detections listed in table A1 of Wilson et al. (2019, hereafter W19). This list contains a total of 154 DA-type111Spectral type DA denotes white dwarfs whose strongest spectral features are the Balmer lines, that typically signify a hydrogen-rich atmosphere, where DAZ denotes weaker lines of metals are present. In contrast, the DZ spectral class are stars with only metal absorption features, where typically the atmosphere is helium dominated. white dwarfs observed in the ultraviolet, 71222Excluding the magnetic object WD 1533–057, as it may be a merger product, and the Si ii lines are likely interstellar (B. Gänsicke 2023, private communication). of which are reported to exhibit photospheric lines of Si ii and are thus DAZ stars. These stars have 14 000 K K, and cooling ages in the range 9–300 Myr.
@@ 4 Wide Binary Fractions and Discussion
There are a handful of companion candidates that are worth discussing in more detail. When the search criteria was eased for the DZ stars, so that their proper motion differences were allowed to be within of a Keplerian orbit, the five candidate binaries thus retrieved for D07 all have , and are thus probable chance alignments. Similarly, two of the three candidate binaries in H18 that resulted from the loosened criteria also have . Only one DZ companion candidate has , and would otherwise be classified as a high-confidence binary: SDSS J091621.36+254028.4 with companion LSPM J0916+2539 (Gaia DR3 687914402017314816). The co-moving companion has an astrometric excess noise of 4.6 mas, a RUWE of 4.8, and is thus an unresolved binary, making the system a hierarchical triple. Nevertheless, it might be argued that the H18 DZ sample has a multiplicity fraction of per cent, which is still drastically lower than for the field, or the DAZ stars. It is noteworthy that the wide companion of the DZ white dwarf SDSS J080740.68+493059.3 is not retrieved in this study, as the latter fails the quality cuts imposed in Section 2.
@@ Acknowledgements
The authors thank the anonymous referee, whose comments helped improve the manuscript. This work presents results from the European Space Agency space mission Gaia, where data are processed by the Gaia Data Processing and Analysis Consortium.

### 19 Mar 2024 v2

# White Dwarf Pollution: One Star or Two?

@@ Abstract
The accretion of tidally disrupted planetary bodies is the current consensus model for the presence of photospheric metals commonly detected in white dwarfs. While most dynamical studies have considered a single star and associated planetary instabilities, several investigations have instead considered the influence of widely-bound stellar companions as potential drivers of white dwarf pollution. This study examines the prevalence of wide binaries among polluted white dwarfs using Gaia Data Release 3 astrometry, where three samples are investigated: 71 DAZ stars with metals detected in the ultraviolet using Hubble Space Telescope, and two groups of DZ stars identified via Sloan Digital Sky Survey spectroscopy, comprised of 116 warmer and 101 cooler sources. Each sample was searched for spatially-resolved, co-moving companions, and compared to the same analysis of thousands of field white dwarfs within overlapping regions of the Gaia Hertzsprung–Russell diagram. The wide binary fraction of the DAZ sample is per cent, and within of the corresponding field. However, the search yields wide binary fractions of less than 1.8 per cent for the two independent DZ star catalogues, which are each distinct from their fields by more than . Both sets of results support that pollution in white dwarfs is not the result of stellar companions, and the delivery of metals to white dwarf surfaces is caused by major planets. The discrepancy between the DAZ and DZ star wide binary fractions cannot be caused by white dwarf spectral evolution, suggesting these two populations may have distinct planetary architectures.
@@
Results to date have found the frequency of atmospheric pollution among single white dwarfs to be indistinct from those in binary systems (Wilson et al., 2019; Zuckerman, 2014) - findings that should be taken with caution as the sample statistics (12 and 38 binaries examined for pollution, respectively) are far from robust. A search for infrared excesses around an unbiased sample of 14 polluted white dwarfs in wide binaries, revealed none to have a significant flux or colour excess, while an independent assessment using Gaia Data Release 2 astrometry further confirmed a duplicity fraction consistent with zero for the 40 known dusty white dwarfs with infrared excesses (Wilson et al., 2019). Previous efforts to identify binary companions or duplicity fractions among white dwarfs have been extremely limited by incomplete or inhomogeneous proper motion and photometric surveys (Farihi et al., 2005; Hollands et al., 2018b). Far superior constraints on the wide binary population of polluted white dwarfs can now be obtained with Gaia.
@@ 2 Sample Selection and Processing
There are three samples of metal-rich white dwarfs utilized to study the fraction of wide binary companions that may dynamically influence their respective planetary systems and pollution characteristics. The first sample is a subset of stars with photospheric metal detections listed in table A1 of Wilson et al. (2019, hereafter W19). This list contains a total of 154 DA-type111Spectral type DA denotes white dwarfs whose strongest spectral features are the Balmer lines, that typically signify a hydrogen-rich atmosphere, where DAZ denotes weaker lines of metals are present. In contrast, the DZ spectral class are stars with only metal absorption features, where typically the atmosphere is helium dominated. white dwarfs observed in the ultraviolet, 71222Excluding the magnetic object WD 1533–057, as it may be a merger product, and the Si ii lines are likely interstellar (B. T. Gänsicke 2023, private communication). of which are reported to exhibit photospheric lines of Si ii and are thus DAZ stars. These stars have 14 000 K K, and cooling ages in the range 9–300 Myr.
@@ 4 Wide Binary Fractions and Discussion
There are a handful of companion candidates that are worth discussing in more detail. When the search criteria was eased for the DZ stars, so that their proper motion differences were allowed to be within of a Keplerian orbit, the five candidate binaries thus retrieved for D07 all have , and are thus probable chance alignments. Similarly, two of the three candidate binaries in H18 that resulted from the loosened criteria also have . Only one DZ companion candidate has , and would otherwise be classified as a high-confidence binary: SDSS J091621.36+254028.4 with companion LSPM J0916+2539 (Gaia DR3 687914402017314816). The co-moving companion has an astrometric excess noise of 4.6 mas, a RUWE of 4.8, and is thus an unresolved binary, making the system a hierarchical triple. Nevertheless, it might be argued that the H18 DZ sample has a multiplicity fraction of per cent, which is still drastically lower than for the field, or the DAZ stars. It is noteworthy that the wide companion of the DZ white dwarf SDSS J080740.68+493059.3 is not retrieved in this study, as the primary fails the quality cuts imposed in Section 2.
@@ Acknowledgements
The authors thank the anonymous referee, whose comments helped improve the manuscript. This work presents results from the European Space Agency space mission Gaia, where data are processed by the Gaia Data Processing and Analysis Consortium. MAH was supported by grant ST/V000853/1 from the Science and Technology Facilities Council (STFC). ST acknowledges support from the Netherlands Research Council NWO (VENI 639.041.645 and VIDI 203.061 grants).